CW-CRDS OF THE OZONE MOLECULE IN THE 5905-5960 cm^{−1} REGION : FIRST OBSERVATION AND ANALYSES OF THE ν_{1}+3ν_{2}+3ν_{3} AND 4ν_{1}+ν_{2}+ν_{3} BANDS

\underline{A. BARBE}, M.-R. DE BACKER-BARILLY, Vl.G. TUYTEREV, Groupe de Spectrométrie Moléculaire et Atmosphérique, U.M.R. CNRS 6089, Université de REIMS, Moulin de la Housse, B.P. 1039, 51687 REIMS cedex 2, FRANCE; A. CAMPARGUE, S. KASSI, Laboratoire de Spectrométrie Physique, U.M.R. CNRS 5588, Université Joseph Fourier B.P. 87, 38402 Saint Martin d’Hères cedex, FRANCE.

This work continues the systematic study of the rovibrational spectra of ozone in the infrared^{a}, newly using the high sensitivity CW-cavity ring down spectroscopy (α_{min} ≈ 3 × 10^{−10} cm^{−1})^{b} ^{c} ^{d} ^{e}. Here, we observe two new A type bands, where the main contributers are ν_{1}+3ν_{2}+3ν_{3} and 4ν_{1}+ν_{2}+ν_{3} centered at 5919.154 and 5947.071 cm^{−1} respectively. We report for each of them the Hamiltonian parameters, the transition moment parameters, the range of observed quantum numbers, the statistics for line position and intensities, the list of ”observed energy levels”, with the comparison with their calculations and we show various examples of agreements between observed and calculated spectra. In addition, a comparison between observed and predicted band center ^{f} will also be discussed.

ahttp://ozone.univ-reims.fr and http://ozone.iao.ru

bM.-R. De Backer-Barilly, A. Barbe, Vl.G. Tyuterev, D. Romanini, B. Moeskop, A. Campargue, J. Mol. Structure, 780 − 781, 225-233, (2006).

cA. Campargue, S. Kassi, D. Romanini, A. Barbe, M.-R. De Backer-Barilly, Vl.G. Tyuterev, J. Mol. Spectrosc., 240, 1-13, (2006).

dA. Barbe, M.-R. De Backer-Barilly, Vl.G. Tyuterev, A. Campargue, D. Romanini, S. Kassi, J. Mol. Spectrosc., under press, (2007).

eA. Barbe, M.-R. De Backer-Barilly, Vl.G. Tyuterev, A. Campargue, D. Romanini, S. Kassi, J. Mol. Spectrosc., submitted, (2007).

fVl.G. Tyuterev, S.A. Tashkun, H. Seghir, A. Barbe, 19^{th} International Conference on High Resolution Molecular Spectroscopy, Praha, (2006), to be published.